OpenStudy (bananas1234):

Which of the following is not a polynomial? z+1 2x^4-y 6 + w y^2-^3 sqrt y + 4

OpenStudy (bananas1234):

6+w? @SolomonZelman

OpenStudy (solomonzelman):

6+w has whole number powers, doesn't it. 6w^0+w^1

OpenStudy (bananas1234):

yes

OpenStudy (solomonzelman):

and 2x^4-y^1 also has whole number powers....

OpenStudy (bananas1234):

yeah

OpenStudy (solomonzelman):

now, do the rest of your options (write the powers)

OpenStudy (bananas1234):

y^2-^3 sqrt y + 4^1 Like that?

OpenStudy (solomonzelman):

but that term sqrt(y) also has a power.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt[3~]{y~} =\sqrt[3~]{y~^{1}}=y^{1/3} }\)

OpenStudy (solomonzelman):

is 1/3 a whole number ?

OpenStudy (bananas1234):

no

OpenStudy (bananas1234):

so the correct answer would be D because 1/3 is not a whole number and the rest have whole numbers in them.

OpenStudy (solomonzelman):

yes

OpenStudy (bananas1234):

Thank you.

OpenStudy (solomonzelman):

the rest have whole number powers in them.

OpenStudy (solomonzelman):

yw